### v2

# ClearDepth: Efficient Stereo Perception of Transparent Objects for Robotic Manipulation

Abstract
Transparent object depth perception remains a major challenge in robotics and logistics due to the limitations of standard 3D sensors in capturing accurate depth on transparent and reflective surfaces. This affects applications relying on depth maps and point clouds, particularly in robotic manipulation. To address this, we propose a vision transformer-based algorithm for stereo depth recovery of transparent objects, enhanced by a novel feature post-fusion module that refines depth estimation using structural image features. To mitigate the high costs of stereo dataset collection, we introduce a physically realistic, domain-adaptive Sim2Real framework for efficient data generation. Our method, accelerated by AI, demonstrates strong Sim2Real generalization, enabling precise depth mapping of transparent objects for robotic applications. Dataset and project details are available at https://sites.google.com/view/cleardepth/.
I INTRODUCTION
Transparent objects, such as glass bottles or cups, are commonly found in domestic service robotics and logistics goods sorting scenarios. Their transparent nature, particularly the complexity of refraction and reflection, poses challenges for imaging and recognition [1]. Accurate handling of these objects is crucial for robots that rely on precise three-dimensional information. Whether it’s accurately grasping glass items in home services or efficiently sorting transparent goods at logistics centers, advanced algorithms and technologies are required for machine vision systems. This includes the development of algorithms capable of handling materials with high reflectivity and refraction, or new sensor technologies to better capture the characteristics of transparent objects. These advancements will enhance the ability of robots to handle transparent objects in various environments, push the application of robotic vision technology into broader fields, and enable robots to perform complex tasks more autonomously. In addressing the depth perception challenges posed by transparent objects, researchers have developed innovative approaches. Initial efforts focused on using deep learning techniques to reconstruct depth information from incomplete depth images captured by RealSense cameras, optimizing for greater accuracy [2, 3]. Subsequent studies explored stereo vision systems [4] and multi-view approaches [5], leveraging neural networks to enhance perception capabilities. Despite these advancements, real-world applications still face challenges, such as inconsistent depth inputs and the complexities of multi-view imaging. A major obstacle in 3D reconstruction of transparent objects is the scarcity of reliable feature points. Innovations addressing this include extracting structural details, such as boundary identification [6], and improving hardware capabilities for more accurate depth imagery [7]. Deep learning has played a pivotal role in learning and interpreting the complex geometric features of these objects, thereby enhancing the precision and reliability of 3D models. For deep learning tasks involving transparent objects, extensive and accurate datasets are crucial. Various techniques have been used for dataset collection, including using markers for object poses [8, 9], substituting transparent objects with opaque ones [10], and manual manipulation in 3D software [11]. However, these methods are labor-intensive and often result in noisy depth maps. To overcome these challenges, simulation engines are increasingly used to generate synthetic data [11, 5], though balancing the realism of ray-tracing with the efficiency of rasterization remains a significant challenge. Our approach includes developing an AI-accelerated tool for generating realistic simulation datasets for transparent object depth recovery. This tool enables direct application of models on actual sensors without post-training, producing instance segmentation, poses, and depth maps. We also introduce a stereo depth recovery network specifically designed for transparent objects. Utilizing a cascaded vision transformer (ViT) backbone, it efficiently extracts contextual structural information. This network outperforms previous CNN and ViT-based models, particularly for high-resolution images and robotic manipulation. Our contributions address the limitations of existing depth estimation methods and datasets for transparent objects. We present a novel AI-enhanced stereo-based simulation dataset generation method, creating a detailed, noise-free dataset. Additionally, our end-to-end stereo depth recovery network operates independently of mask priors and includes a post-processing structure that enhances structural details. This stereo imaging algorithm can be directly integrated into existing robotic grasping pipelines. The whole pipeline is shown in Fig. LABEL:fig.cleardepth_overview.
II Related Work
II-A Perception of Transparent Objects
Robotic perception of transparent objects remains challenging due to their low contrast and complex light interactions, which affect sensor accuracy in determining position and shape. Traditional RGB and RGB-D cameras struggle with these objects, as they rely on intensity data and overlook optical properties. To address this, research has explored polarized cameras, which reduce reflections and enhance contrast [7, 12]. However, their high cost limits widespread adoption. Alternative approaches include CNN-transformer-based models for tracking [13], Sim2Real techniques leveraging synthetic datasets [14], and alpha-matting methods for transparent object segmentation [15]. For robotic manipulation and pose estimation, multi-task perception models have been introduced [8, 11, 16]. Depth recovery remains particularly challenging due to light refraction and reflection. Methods such as NeRF and volumetric rendering aid surface reconstruction [17, 18, 19], while stereo and multi-view techniques improve depth estimation [4, 20, 9]. These approaches leverage various sensors, including RGB-D, stereo vision, and multi-view systems, to enhance transparent object perception [3, 2, 8, 11, 16, 9, 4, 5, 18, 7]. Advances in deep learning and sensor technologies continue to drive improvements in accuracy and reliability.
II-B Deep Learning-based Stereo Depth Recovery
Deep learning-based stereo matching methods have recently outperformed traditional approaches, with 2D convolutional models [21, 22] offering simplicity and efficiency. These models achieve high accuracy even on limited computational resources, making them suitable for engineering applications, though they still require improvements in accuracy and robustness due to 3D cost space constraints. 3D convolutional networks [23, 24] provide better interpretability and higher disparity map accuracy but require optimization due to their computational demands. RAFT-Stereo [25], an extension of RAFT [26], applies optical flow techniques to stereo matching, improving generalization and robustness with its lightweight Gated Recurrent Unit (GRU) Network module, but struggles with global context extraction due to its CNN architecture. STTR [27], inspired by SuperGlue [28], uses transformers with positional embedding and attention mechanisms for binocular dense matching, producing disparity and depth maps. However, these methods are computationally intensive and slow in inference, limiting their suitability for high-resolution images and downstream robotic tasks.
II-C Transparent Object Datasets
Recent works [29, 30] show that real-world datasets degrade model performance due to label noise, while synthetic data with precise labels, enhance model performance. However, synthetic datasets across different data domains remain scarce in the open-source community. Most recent vision tasks leverage ray-tracing renderers, minimizing the sim2real gap, making data domain gaps the primary factor affecting model performance. Therefore, continuously expanding synthetic datasets in the open-source community is crucial. Existing synthetic datasets, such as  [10, 31, 9, 32, 33, 16], typically feature transparent objects on desktops. However, these datasets lack the complexity necessary for generalization to real-world scenarios like kitchens, bedrooms, and offices, where service and humanoid robots operate. Moreover, these datasets often require extensive pre- or post-processing, such as segmentation or background reconstruction, which is impractical for end-to-end algorithms crucial to embodied intelligence applications. Synthetic datasets with HDRI backgrounds, such as [5], also face challenges. HDRI backgrounds lack depth value labels, leading to confusion and poor generalization, especially in zero-shot scenarios for service robots. Simulation data for Realsense cameras, like  [16, 33], involve complex rendering processes that reduce efficiency. The process includes rendering the RGB image, simulating IR projection with spotlights, and generating noisy depth maps, which is both complex and time-consuming. In designing our dataset and generation process, we address the limitations of existing datasets while anticipating the needs of future embodied intelligence algorithms. Our dataset includes indoor scenes typical for service robots, enriched with background depth values for transparent objects. It features both container-type and cosmetic transparent objects, carefully designed to overcome current deficiencies and ensure future scalability.
III Method
III-A Network Overview
Figure 2: Our stereo depth recovery network for transparent objects. Feature extraction is performed on left and right images, with additional processing by a shared weighted context encoder for the left image. A correlation pyramid, created by merging the feature maps, is refined through a GRU loop, enhancing structural information for depth recovery. The output is a refined disparity map. The feature encoder is used for computing visual similarity for left and right image, the context encoder extracts features for disparity refinement in the GRU.
Transparent objects refract background textures, making structural features more critical than texture features for imaging and perception. [34] shows CNNs excel at texture recognition, while ViTs are better at shape recognition. Traditional ViTs downsample inputs and use learnable upsampling, which, while effective, is computationally costly and weak in fine-grained feature extraction. Models like SegFormer [35] and DinoV2 [36] enhance ViTs with cascaded structures, improving depth estimation and segmentation through multi-scale features. To better capture structural features for transparent objects, we adopt an efficient cascaded ViT backbone. Traditional stereo networks rely on dot product similarity, which fails for transparent objects due to background refraction. Instead, we introduce structural feature priors into the GRU loop via a lightweight post-fusion module, enhancing performance without the overhead of cross-attention. To ensure adaptability to robotic tasks, we keep the model architecture lightweight to avoid potential engineering challenges. Our network is illustrated in Fig. 2.
III-A1 Cascaded Vision Transformer Backbone.
Our backbone begins with overlap patch embedding for initial tokenization, preserving local features. Tokens pass through four transformer blocks, generating feature maps at $\frac{1}{4}$, $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$ scales. To optimize computational efficiency, the model incorporates efficient self-attention, which significantly reduces the computational burden from $O(N^{2})$ to $O(\frac{N^{2}}{R})$. This reduction is achieved by first reshaping the input sequence from $N\cdot C$ to $\frac{N}{R}\times(C\cdot R)$ by 2d convolutional layer with the stride $8,4,2,1$ for different ViT blocks, as detailed in Equ. 1, and then adjusting the sequence dimensions back to $C$ channel through linear layers, as described in Equ. 2. K denotes the sequence in the ViT block optimized for lower computational complexity.
$\hat{K}=Reshape(\frac{N}{R},C\cdot R)(K)$ (1)
$K=Linear(C\cdot R,C)(\hat{K})$ (2)
Additionally, the Mix-FFN module in the architecture addresses the challenge of performance degradation due to the interpolation of positional embeddings in the original ViT structure, especially when dealing with varying input image sizes, here by substituting positional embeddings with learnable depth-wise convolutions. The equation is as 3.
$\mathbf{x}_{\text{out}}=\text{MLP}(\text{GELU}(\text{Conv}_{3\times 3}(\text{MLP}(\mathbf{x}_{\text{in}}))))+\mathbf{x}_{\text{in}}$ (3)
Then, we concatenate multi-scale feature maps from different ViT blocks by upsampling them to a unified scale of $\frac{1}{4}$. This combined feature map undergoes further refinement through a precise $1\cdot 1$ convolution, facilitating optimal dimension adjustment.
III-A2 Structural Feature Post-Fusion
In this work, we propose a modified GRU-based architecture to optimize disparity maps in a coarse-to-fine manner. Our Post-Fusion modification is particularly introduced to enhance the processing of transparent objects. We made this modification because, during our depth estimation experiments on transparent objects, we observed that unlike regular objects, depth estimation for transparent objects particularly requires precise structural information. Additionally, for transparent objects, due to the refractive nature of their background texture features, the feature similarity information obtained through dot product is not suitable for reconstructing transparent objects. Therefore, it makes sense to introduce the structural information of the image itself during the GRU iterations. This approach ensures that the structural information extracted at different resolutions remains consistent throughout the iterative process.
The core update equations in our model are defined as follows:
$x_{k}=$ $\;[\mathbf{C}_{k},\mathbf{d}_{k},\mathbf{c}_{k},\mathbf{c}_{r},\mathbf{c}_{h}]$ (4)
$z_{k}=$ $\;\sigma(\text{Conv}([h_{k-1},x_{k}],W_{z})+c_{k}),$ (5)
$r_{k}=$ $\;\sigma(\text{Conv}([h_{k-1},x_{k}],W_{r})+c_{r}),$ (6)
$\tilde{h}_{k}=$ $\,\tanh(\text{Conv}([r_{k}\odot h_{k-1},x_{k}],W_{h})+c_{h}),$ (7)
$h_{k}=$ $\;(1-z_{k})\odot h_{k-1}+z_{k}\odot\tilde{h}_{k},$ (8)
Here, $x_{k}$ is a concatenation of several feature maps, including the correlation $\mathbf{C}_{k}$, the current disparity $\mathbf{d}_{k}$, and structural context feature maps $\mathbf{c}_{k}$, $\mathbf{c}_{r}$, and $\mathbf{c}_{h}$. Specifically, $\mathbf{c}_{k}$, $\mathbf{c}_{r}$, and $\mathbf{c}_{h}$ represent structural features derived from the left image. These features are incorporated as residuals into the GRU loop, allowing for enhanced participation of structural information during the disparity map refinement process. $z,r,h$ represent the state information of the update gate, reset gate, and hidden gate in a GRU.
Figure 3: We utilize hardware-accelerated ray tracing, pretrained OptiX AI denoiser, and pretrained AI-driven super-resolution. This process accelerates the dataset generation process.
Figure 4: The left image displays rendered models, with the first two rows featuring transparent containers such as vases, bowls, and cups, while the last two rows showcase cosmetic products made of transparent glass and plastic. The right image shows the real objects used for Sim2Real testing.
Then, Our approach decode GRUs at each resolutions to obtain multi-scale disparity updates for coarse to fine gradual optimization:
$\triangle\mathbf{d}_{k,\frac{1}{32}}$ $=\text{Decoder}(h_{k,\frac{1}{32}}),$ (9)
$\triangle\mathbf{d}_{k,\frac{1}{16}}$ $=\text{Decoder}(h_{k,\frac{1}{16}}+\text{Interp}(\triangle\mathbf{d}_{k,\frac{1}{32}})),$ (10)
$\triangle\mathbf{d}_{k,\frac{1}{8}}$ $=\text{Decoder}(h_{k,\frac{1}{8}}+\text{Interp}(\triangle\mathbf{d}_{k,\frac{1}{16}})),$ (11)
where Decoder consist of two convolutional layers and Interp is bilinear interpolation scaled up by a factor of two. Finally, the updated disparity is calculated as:
$\mathbf{d}_{k+1}=$ $\;\mathbf{d}_{k}+\triangle\mathbf{d}_{k}$ (12)
In summary, to address the challenges of transparent objects, we selected an appropriate image feature extractor. Additionally, considering the unique difficulties of transparent objects and the need for efficient models in robotics, we designed a structural feature post-fusion architecture. Every detail of our network structure is tailored to the characteristics of transparent object scenarios.
In the comparative experiments section, the visual results demonstrate that our model substantially enhances the stereo imaging of transparent objects.
III-B Synthetic Dataset Generation
Figure 5: Sample images from SynClearDepth dataset, depicting transparent objects randomly placed in indoor scenes (bathroom, dining room, kitchen, living room) under various lighting conditions. The objects, including cosmetic packaging and glass containers, are randomly dropped onto tables using 3D bounding boxes as collision bodies.
To enhance the efficiency of synthetic dataset generation, we utilized the AI denoiser provided by OptiX [37] during rendering and adopted open-source pretrained deep learning super-resolution [38] as rendering output optimization strategies, reducing the average generation time per set (stereo RGB, depth, masks, and object-camera poses) from $12.77$ to $4.40$ seconds. Since these techniques are widely used in computer graphics and do not alter the core data distribution, we omit further analysis. The dataset generation process is illustrated in Fig. 3. Our SynClearDepth dataset includes 16 selected objects: 10 common transparent containers and 6 glass-material products (Fig. 4). To ensure depth labels for both objects and backgrounds, we combined object models with indoor scenes, including 6 bathrooms, 3 dining rooms, 5 kitchens, and 6 living rooms. This resulted in 14,091 image sets, each containing left and right RGB images, ground truth depth, instance segmentation, and object/camera poses (Fig. 5). We applied domain randomization to object types, quantities, poses, lighting, and camera angles. This dataset is designed to support robotic perception and manipulation in service robot applications, particularly for handling transparent objects in household environments.
IV Experiments
IV-A Technical Specifications
Our network is firstly pre-trained on CREStereo dataset [39] and Scene Flow dataset [40], and then fine-tuned on our proposed SynClearDepth dataset for transparent object stereo imaging. Our model is trained on 1 block of NVIDIA RTX A6000 with batch size 8 and the whole training lasts for 300,000 steps. We use AdamW [41] as optimizer, the learning rate is set to 0.0002, updated with a warm-up mechanism and used one-cycle learning rate scheduler. The final learning rate when training finished is 0.0001. The input size of the model is resized to $360\times 720$. Fine-tune for transparent objects takes the same training parameters as pretraining on the opaque dataset.
IV-B Evaluation Metrics
  1. 1.
    AvgErr (Average Error): Represents the average disparity error across all pixels, indicating the general accuracy of the disparity map.
  2. 2.
    RMS (Root Mean Square Error): Measures the square root of the average squared disparity error, reflecting the overall deviation from the ground truth.
  3. 3.
    Bad 0.5 (%), Bad 1.0 (%), Bad 2.0 (%), Bad 4.0 (%): These metrics indicate the percentage of pixels where the disparity error exceeds 0.5, 1.0, 2.0, and 4.0 pixels, respectively, highlighting the proportion of significant errors in the disparity map.
Together, these metrics provide a comprehensive assessment of stereo matching performance, balancing both overall accuracy and the frequency of large errors.
Methods AvgErr RMS bad 0.5 (%) bad 1.0 (%) bad 2.0 (%) bad 4.0 (%)
RAFT-Stereo[25] 1.27 8.41 27.7 9.37 4.14 2.75
CREStereo[39] 1.15 7.70 28.0 8.25 3.71 2.04
ClearDepth 1.33 8.68 25.30 7.39 3.48 2.00
TABLE I: Quantitative results on Middleburry Stereo Evaluation Benchmark [42]. All metrics have been calculated using undisclosed weighting factors. The outcomes unequivocally demonstrate that our technique substantially outperforms the baseline method.
Methods AvgErr RMS bad 0.5 (%) bad 1.0 (%) bad 2.0 (%) bad 4.0 (%)
IGEV-Stereo[43] 2.077 5.5301 58.9743 36.1270 19.967 10.777
DLNR[44] 3.097 8.4269 28.1088 21.8442 16.481 11.9046
Selective-IGEV[45] 1.273 4.3365 34.8229 17.6288 9.561 5.8707
RAFT-Stereo[25] 2.245 8.8016 29.7356 17.4521 10.835 6.2107
ClearDepth 2.138 8.7282 24.7329 16.3178 9.8459 5.7600
TABLE II: Quantitative results on transparent object dataset compared with stereo SOTA methods. Visualization results shown in Fig. 7. Our efficient stereo depth estimation model outperforms other SOTA models in transparent object dataset whtile maintaining lower computational cost. Results show that our model achieves strong performance in average numerical error and outperforms all others in the proportion of error pixels.
IV-C Qualitative and Quantitative Studies for Stereo Depth Estimation
IV-C1 Quantitative Analysis on Middlebury Dataset
The Middlebury 2014 dataset comprises 23 pairs of images designate for training and validation purposes. We refine our model over these 23 pairs, conducting fine-tuning across 4,000 iterations with an image resolution of $384\times 1024$. Benchmark against standard baseline approaches RAFT-Stereo and CREStereo using various stereo evaluation metrics further underscores the efficacy of our approach, as outlined in Tab. I. More comparison results with other methods can be found at [42].
IV-C2 Quantitative Analysis on KITTI Dataset
We fine-tune our pre-trained model using the KITTI 2015 training set across 5,000 steps, employing image crops sized at $320\times 1000$. The learning rate is established at 0.00001, with the batch size held at 3. In terms of GRU updates, we perform 22 iterations during training, adjusting to 32 iterations for testing. Given that the labels and metrics of the KITTI dataset cannot fully reflect imaging quality [29], we conducted a qualitative analysis on the KITTI dataset. Fig. 6 demonstrates a competitive comparison focused on detail recovery, our method shows exceptional proficiency in reconstructing depth details of foreground objects, significantly outstripping alternative approaches by a substantial margin.
(a) Left image
(b) RAFT-Stereo[25]
(c) StereoBase[46]
(d) Ours
Figure 6: Visual comparisons on KITTI 2015 with SOTA StereoBase and baseline RAFT-Stereo. Our method is more robust to overall scene details.
IV-C3 Evaluation on Transparent Object Dataset
Figure 7: The visualization results of our transparent object stereo depth reconstruction method compare with other SOTA stereo depth estimation methods by fine-tuning on SynClearDepth dataset.
To validate our model and dataset for transparent object depth recovery in stereo vision, we fine-tuned our pre-trained model on the SynClearDepth dataset using the same training parameters as pre-training. We also fine-tuned RAFT-Stereo [25] and other SOTA models from the Middlebury benchmark on SynClearDepth for comparison. This analysis highlights our model’s improvements in stereo-based depth perception for transparent objects. Tab. II presents quantitative results, while Fig. 7 visualizes stereo imaging performance. AvgErr (Average Error) and RMS (Root Mean Square Error) measure numerical error, while Bad 0.5 (%), Bad 1.0 (%), Bad 2.0 (%), Bad 4.0 (%) reflect relative error. Results show our model achieves strong performance in numerical error and outperforms all others in relative error. Our model is more efficient than others, achieving comparable performance without the high computational cost of cross-attention or multi-model ensembles. This is due to innovations in the image encoder, making our approach more suitable for robotics. Additionally, our TensorRT implementation enables real-time inference at 50 FPS on consumer GPUs, whereas other models, due to their complex designs, are impractical for deployment. We perform qualitative analysis on real-world images. Results in supplementary video show that leveraging a physically realistic renderer enables strong generalization in real world, with performance consistent across domain-shifted test sets. In this work, we adopt a stereo-based approach instead of the Realsense-based imaging methods [33]. Realsense cameras, as consumer-grade devices, do not meet industrial standards. Moreover, their open-source datasets are limited in both diversity and scale compared to stereo datasets, restricting future extensions. Additionally, Realsense IR projection penetrates transparent objects, leading to visual information loss, which we avoid to ensure robustness.
Methods AvgErr RMS bad 0.5 (%) bad 1.0 (%) bad 2.0 (%) bad 4.0 (%)
w/o Fusion 6.90 15.48 43.34 29.63 21.52 16.62
Feature Fusion 2.64 8.59 27.23 16.87 11.28 7.72
TABLE III: Ablation study for the feature post-fusion module in clearDepth with 100,000 steps on SynClearDepth dataset.
IV-C4 Ablation Study of Feature Post-Fusion Module
Object Number 1 2 3 4 5
success rate 98% 92% 86% 78% 76%
TABLE IV: Grasping success rate with varying number of transparent objects.
To evaluate the impact of our feature post-fusion module, we conducted ablation studies on the SynClearDepth dataset. We compared networks with and without the module, as shown in Tab. III. Results indicate a substantial performance boost, especially in handling complex transparency and light refraction, highlighting its effectiveness in enhancing depth estimation and object recognition. Each study was trained for 100,000 steps.
IV-D Robotic Grasping Experiments
We employ the ZED 2 stereo camera with our network, integrated with a robotic arm and five-fingered hand, for transparent object grasping. Following grasping pipeline from  [47], we conducted 150 trials per scenario, varying object counts from 1 to 5, achieving an 86.2% average success rate (Tab. IV). Our method seamlessly integrates into existing grasping frameworks without extra intermediate processing or multi-view capture, ensuring efficiency. Additionally, our fast data generation enables rapid dataset expansion. Results show that our depth recovery network substantially enhances transparent object grasping.
V Conclusion and Future Work
In this work, we present a complete visual perception framework for transparent object manipulation in service robotics scenarios, spanning synthetic data generation, stereo depth estimation, and real-world robotic validation. We propose an efficient real-time stereo depth recovery network that combines a cascaded vision transformer backbone with a structural feature post-fusion module, enabling fine-grained structural perception and accurate depth recovery of transparent objects without relying on mask priors. To address the data scarcity challenge in transparent object perception, we construct SynClearDepth, a high-quality simulation dataset containing diverse household environments and realistic object placements. It provides accurate RGB, depth maps, instance masks, and pose annotations, substantially enhancing model generalization in real-world scenarios. We validate our model through extensive comparisons on public and proprietary datasets, along with ablation studies. Experimental results demonstrate that our approach outperforms existing methods on both public and proprietary benchmarks, particularly in structure-aware and boundary-level depth estimation. Results demonstrate its robustness, accuracy and efficiency, supporting transparent object manipulation in robotics. Furthermore, real-world robotic grasping experiments show that our method can be seamlessly integrated into grasping pipelines without requiring multi-view capture or additional pre-processing, and achieves stable and precise manipulation of transparent objects. These results highlight the practicality and applicability of stereo-based transparent object depth estimation in real-world robotic tasks.
Expanding stereo datasets across domains is an ongoing effort. Recent research on visual foundation models highlights the critical role of high-quality synthetic data, yet the open-source community still lacks sufficient quantity and diversity. Future work will focus on expending stereo data for robotic grasping, enabling commercial deployment in logistics sorting for transparent and reflective objects like cosmetics.